OpenStudy (shaunamariexoxo):

A nationwide club begins a chapter near you. You research the memembership of the club over the past few decades. The table shows your data what is the equation for a line of best fit? How many members would you expecf to be in the year 2019? The table looks like: YEAR: 1980 MEMBERSHIP:5,7000 YEAR: 1985 MS: 7,200 YEAR: 1990 MS: 8,600 YEAR: 1995 MS: 10,100 YEAR: 2000 MS: 11,500 YEAR 2005 MS: 12,500

OpenStudy (retireed):

So how would you work this, I have an idea but want you to tell me what you think.

OpenStudy (shaunamariexoxo):

I have no idea I've been trying to figure it out forever. I tried adding all the memberships but I don't think that's right

OpenStudy (retireed):

I plot the data point for the every 5 year period I noted the delta between each 5 year data point and it peaked at 1995 and started trend down I just plotted the future data trending down slowly. Does this make sense?

OpenStudy (retireed):

MEMBERSHIP Delta YEAR: 1980: 5,700 YEAR: 1985 MS: 7,200 1500 YEAR: 1990 MS: 8,600 1400 YEAR: 1995 MS: 10,100 1500 YEAR: 2000 MS: 11,500 1400 YEAR 2005 MS: 12,500 1000 Is the last membership number typed correct? Is it really 12,500?

OpenStudy (retireed):

If the last membership number was 12,900 I would see better trend for the analysis.

OpenStudy (shaunamariexoxo):

Yes it's 12,500 on my paper :(

likeabossssssss (likeabossssssss):

\(\bbox[5pt,midnightblue]{\LARGE\color{yellow}{\star ^\bigstar Welcome~to~~os~ ^\bigstar \star}}\)

OpenStudy (retireed):

YEAR: 1980: 5,700 YEAR: 1985 MS: 7,200 1500 = 7200 - 5700 YEAR: 1990 MS: 8,600 1400 = 8600 - 7200 YEAR: 1995 MS: 10,100 1500 YEAR: 2000 MS: 11,500 1400 YEAR 2005 MS: 12,500 1000 the trend seems to be declining YEAR: 2010 MS: 13,100 600 I kept a declining trend YEAR: 2015 MS: 13,400 300 YEAR: 2020 MS: 13,450 50 So in my estimate at year 2019 the membership will be _____________? Your guess is as goo as mine.